1 mol of MnO4- will oxidise x mol of ferric oxalate in acidic medium, x is---------.Answer: 5/6

Grade:12th pass

1 Answers

Vikas TU
7 years ago
MnO4- will oxide to MnO2 thus its v.f  would be 5 and Fe3+ would remain fe3+ a dn no. of atoms of Fe3+ is 2 therefore,
moles * valence factor = moles * valence factor
1 * 5 = x * 3*2
 
x = 5/6